# What is the modular Fourier transform?

### v2

@@ Abstract.
The discrete Fourier transform can be viewed as an isomorphism of the group algebra of a cyclic group to its decomposition in terms of irreducible representations when the characteristic does not divide the order of the group. We describe the DFT when $p|N$ by factoring $x^{N}-1$ over finite fields. For the symmetric group, the Fourier transform is implemented using representations, Specht modules $S^{\lambda}$ for partitions $\lambda$, and Maschke’s theorem. In the modular case, we construct the Peirce decomposition using central primitive orthogonal idempotents $e_{i}$ in SageMath, yielding a change-of-basis matrix which generalizes the DFT.
@@ 1. Introduction
The Fourier transform is an old and powerful tool used in variety of applications, from signal processing to fundamental physics. It is central in mathematics. At its core, it is a change of basis that allows us to view a signal changing in time in terms of static frequencies. This article aims to build intuition for an extension of the usual Fourier transform which takes place over the field of real or complex numbers to a more general field of potentially positive characteristic. Further, we implement the so called “modular Fourier transform” in SageMath for finite abelian groups and for the symmetric group $S_{m}$ for $m\leq 6$. This requires computing the dimensions of the simple modules $D_{p}^{\lambda}$ of the symmetric group. We do so by computing the rank of the Gram matrix over $\mathbb{F}_{p}$. The modular Fourier transform for a field $K$ with $\text{char}(K)>0$ uses the isomorphism from the group algebra $K\left[G\right]$ to its block decomposition $\bigoplus_{i}B_{i}$, where $B_{i}=K\left[G\right]e_{i}$ and $1=\sum_{i}e_{i}$ is a decomposition of the identity into primitive orthogonal idempotents. This generalizes the usual Fourier transform which uses Mashke’s theorem to decompose $K\left[G\right]=\bigoplus_{i}End(V_{i})$ for irreducible representations $V_{i}$ of $G$ so long as $\text{char}(K)\nmid|G|$.
@@ Acknowledgements
I would like to extend my sincere thanks to Dave Benson, Ezra Brown, Yves de Cornulier, Jill Pipher, Sridhar Ramesh, Steve Rosenberg, Travis Scrimshaw, Jared Weinstein, and Mark Wildon for helpful comments and conversation.
@@ 2. The Discrete Fourier Transform
The DFT generalizes readily to the realm of quantum computing. We view a collection of $n$ qubits $c_{0}|0\rangle+c_{1}|1\rangle\in\mathbb{C}^{2}$ which are concatenated in space to correspond to the Hilbert space $\mathcal{H}\cong\left(\mathbb{C}|0\rangle+\mathbb{C}|1\rangle\right)^{n}\cong\mathbb{C}^{2^{n}}$. Then the Fourier transform is a unitary matrix $U:\mathcal{H}\rightarrow\mathcal{H}$, where we label the eigenstates using binary notation, e.g. for $n=3$, $|0\rangle\otimes|1\rangle\otimes|1\rangle=|011\rangle$ can be written as $|3\rangle$ in the familiar decimal, and we interpret the label modulo $N=2^{n}$. We can then define maps such as $|x\rangle\mapsto|a^{x}(mod\,N)\rangle$ which is useful in applications like Shor’s algorithm, where we use the inverse Fourier transform to extract the periodicity of that modular exponentiation viewed as a function $\mathbb{Z}/N\mathbb{Z}\rightarrow\mathbb{Z}/N\mathbb{Z}$.
@@
This technique is referred to as “Fourier sampling”. It is important that the base field, $\mathbb{C}$, is of characteristic zero. This means that we can always invert the order of the cyclic group, $N$, to define the inverse Fourier transform. However, we have thus far suppressed the representation theory and harmonic analysis. Note that $\mathbb{Z}/N\mathbb{Z}$ is an abelian group, so its representations are one dimensional. So we’re looking at maps $\chi:\mathbb{Z}/N\mathbb{Z}\rightarrow\mathbb{C}$. This means that $\chi(N)=\chi(1)^{N}=1$, which means $\chi(1)$ is an $N^{th}$ root of unity. Since these are one dimensional, they are irreducible. Thus, there are $N$ rep’ns given by $\chi_{k}(n)=\exp(\frac{-2\pi ikn}{N})$.
@@
Note that $\chi_{k}$ are functions on the group $\mathbb{Z}/N\mathbb{Z}$. Thus, for each group element $n\in\{0,\ldots,N-1\}$ we have a complex number $\chi_{k}(n)\in\mathbb{C}$. Recall that the space of class functions on a group form a vector space. Since the group is abelian, each conjugacy class has one element, so $\chi_{k}$ may be viewed as a vector in $V^{*}\cong\mathbb{C}^{N}$. Separately, we have the group alegbra $V=\mathbb{C}\left[\mathbb{Z}/N\mathbb{Z}\right]\cong\mathbb{C}^{N}$ as a vector space. We can multiply these “vectors” by viewing group elements as a basis and extending the multiplication as we do with polynomials. Label the irreducible representations $V_{k}\cong\mathbb{C}$. Each element $n\in\mathbb{Z}/N\mathbb{Z}$ may be viewed as an endomorphism $n:V\rightarrow V$ by multiplication. That is, if $s=\sum_{j}c_{j}j\in V$ then $n\cdot s=\sum_{k}c_{j}(n\cdot j)$ where the product $n\cdot j=n+j\mod N$ is in $\mathbb{Z}/N\mathbb{Z}$. Therefore we have a linear action of $\mathbb{Z}/N\mathbb{Z}$ on $V$, which is invertible since $-n$ is the multiplicative inverse of $n$. Thus, each $n\in GL(V)$, so this is a representation of the group. $V$ can be decomposed into irreducible representations $V\bigoplus_{k}V_{k}$ and the action of $n$ diagonalized as $n_{k}\in End(V_{k})$. We thus have the decomposition $V=\mathbb{C}\left[\mathbb{Z}/N\mathbb{Z}\right]\cong\bigoplus_{k}End(V_{k})\cong\bigoplus_{k}\mathbb{C}\chi_{k}=V^{*}$.
@@
So how does the Fourier transform look in this interpretation? Well, consider the element $x=\left(x_{1},\ldots,x_{N}\right)\in\mathbb{C}^{N}$. We map it to $\hat{x}_{k}=\sum_{n=0}^{N-1}x_{n}\exp\left(\frac{-2\pi ikn}{N}\right)$.
@@
Let’s rewrite this formula a bit. $\hat{x}_{k}=\sum_{n=0}^{N-1}x_{n}\exp\left(\frac{-2\pi k}{N}\right)^{n}=\sum_{n=0}^{N-1}x_{n}\chi_{k}(n)$. We see that our “Fourier coefficients” are just sums of the components of $x$ against the values of the irreducible characters. We obtain $N$ coefficients, and this is exactly the discrete Fourier transform for the group $\mathbb{Z}/N\mathbb{Z}$. This formula has the nice property of being invertible, with inverse given by $x_{n}=\frac{1}{N}\sum_{k=0}^{N-1}\hat{x}_{k}\chi_{k}(-n)$, whose proof uses the orthogonality of characters.
@@
This is a bit of a special case, in which the number of representations equals the order of the group, i.e. every representation is degree 1. This only occurs for abelian groups. For nonabelian groups, we need to use matrices.
@@ 3. Fourier Transform for Non-abelian Groups
Let $G$ be a finite group and let $\hat{G}$ denote the set of irreducible representations of $G$, also a finite set. Over a field of characteristic zero, we have Maschke’s theorem:
@@
This theorem tells us that as a $K\left[G\right]$-module (multiply polynomials) we may decompose $K\left[G\right]$ into a direct sum of simple $K\left[G\right]$-modules. That is, $K\left[G\right]=\bigoplus_{i}M_{i}$. It turns out these simple $M_{i}$ are just the endomorphism algebras of irreducible representations, i.e. $K\left[G\right]=\bigoplus_{i}End(V_{i})$. Thus, an element of the group algebra may be viewed as a “block diagonal” matrix acting on itself, with each block corresponding to an irreducible representation.
@@
where $f(a)\in K$ and each $\rho(a)$ is a $d_{\rho}\times d_{\rho}$ matrix. This formula is invertible yielding:
$f(a)=\frac{1}{|G|}\sum_{g\in\hat{G}}d_{\rho}\text{Tr}(\rho(a^{-1})\hat{f}(\rho))$
We notice a few things. It is clear we require $\text{char}(K)\nmid|G|$, both in order to comply with the hypothesis of Mashke’s Theorem, and to avoid division by zero. We need to know the dimension numbers of irreducible representations, $d_{\rho}$. We are dealing with matrices, so we are taking the matrix trace now in the sum.
@@ 4. The Symmetric Group over $\mathbb{C}$
For the symmetric group over $\mathbb{C}$, Maschke’s theorem applies. The irreducible representations are labeled by Young tableaux $\lambda$ corresponding to partitions of $m$. For each $\lambda$, we obtain a simple modules $V^{\lambda}$ with basis given by
@@
where $Q_{T}$ is the subgroup of permutations, preserving (as sets) all columns of $T$ and $\epsilon(\sigma)$ is the sign of the permutation $\sigma$. The Specht module of the partition $\lambda$ is the module generated by the elements $E_{T}$ as $T$ runs through all tableaux of shape $\lambda$. The dimension of the Specht module $V^{\lambda}$ is the number of standard Young tableaux of shape $\lambda$. It is given by the hook length formula. Over fields of characteristic 0 the Specht modules are irreducible, and form a complete set of irreducible representations of the symmetric group.
We have everything we need to compute the non-abelian Fourier transform, which is implemented in [FTSG] for the symmetric group (though it generalizes to any finite group given a labeling for the representations).
@@
There is a well-known isomorphism of categories between $K\left[G\right]$-modules are representations of the group $G$ over the field $K$. The “regular representation” where $k\left[G\right]$ acts on itself by multiplication decomposes into a sum of irreducible representations. When the characteristic of the field $p$ does not divide the order of the group $|G|$ we can use Maschke’s theorem as stated above. The key element of the proof is the construction of an “average” or invariant inner product, $\langle x,y\rangle_{\text{inv}}=\frac{1}{|G|}\sum_{g\in G}\langle g\cdot x,g\cdot y\rangle$. Note that typically we require $K$ be algebraically closed, since otherwise a more complicated statement holds: $K\left[G\right]$ is a product of matrix algebras over division rings over $K$. In the case of finite fields, any division ring over $F_{p}$ must be a field by Wedderburn’s little theorem, and the only finite fields of characteristic $p$ are the extensions $F_{q}$ where $q=p^{\nu}$.
When $p\mid|G|$, the situation is more complicated and is the subject of modular representation theory. In our case, we are primarily concerned with cyclic groups and the symmetric group $S_{m}$. In general, we no longer have Maschke’s theorem giving us a decomposition of the group algebra into simple modules. Instead, we have a decomposition into “blocks”:
@@
where $\sum_{i}e_{i}=1$ is a decomposition of the identity into central primitive orthogonal idempotents. Far from being abstract placeholders, in certain cases these idempotents are readily computable.
@@ 6. The Discrete Fourier Transform over a Ring
We implement this procedure in SageMath code. [SAGE] [DFTCG]
@@
When $R=\mathbb{F}_{p}$, note that $\mathbb{F}_{p}^{*}\cong\mathbb{Z}/(N-1)\mathbb{Z}$, that is the multiplicative group of a finite field is cyclic. By Fermat’s little theorem, $\alpha^{p-1}=1$. Thus, we require $N|p-1$. This implies that $p>N+1$, which means that the characteristic we choose must be bigger than the size of the list we are transforming. This is quite a constraint!
@@
#find an n-th root in finite field of size p
def primitive_root(K,n):
for a in K:
if a**n == 1 and all(a**k != 1 for k in range(1,n)):
return a
@@ 7. Cyclic Groups over $\mathbb{F}_{p}$
The question of how to classify all finite-dimensional representations of $\mathbb{Z}/N\mathbb{Z}$ over an arbitrary field $F$ can be studied using the structure theorem for finitely-generated modules over a principal ideal domain, in this case $F\left[x\right]$. The structure theorem asserts that any finitely-generated module is uniquely a finite direct sum of modules of the form $F\left[x\right]/p(x)^{r}$ where $p\in F\left[x\right]$ is irreducible and $r$ is a non-negative integer.
If $T$ is an operator acting on $F^{k}$ for some $n$, then $F^{k}$ becomes a finitely-generated module over $F\left[x\right]$ with $x$ acting by $T$. $T$ gives a representation of the cyclic group $\mathbb{Z}/N\mathbb{Z}$ if and only if $T^{N}=1$, in which case the summands $F\left[x\right]/q(x)^{r}$ in the decomposition of $F^{k}$ must have the property that $q(x)^{r}|x^{N}-1$.

### v3

# The Modular DFT of the Symmetric Group

@@ Abstract.
We describe the discrete Fourier transform (DFT) for a cyclic group when $p|N$ by factoring $x^{N}-1$ over finite fields and constructing the Fourier transform and its inverse using Bézout’s identity for polynomials. For the symmetric group, in the modular case when $p|n!$ we construct the Peirce decomposition using central primitive orthogonal idempotents, yielding a change-of-basis matrix which generalizes the DFT. We compute the unitary DFT for the symmetric group over number fields containing sufficiently many square roots. For $n=3$, we compute the Galois group of the splitting field of the characteristic polynomial. All constructions are implemented in SageMath.
@@ 1. Introduction
The Fourier transform is an old and powerful tool used in variety of applications, from signal processing to fundamental physics. It is central in mathematics. At its core, it is a change of basis that allows us to view a signal changing in time in terms of static frequencies. This article aims to build intuition for an extension of the usual Fourier transform that takes place over the field of real or complex numbers to a more general field of potentially positive characteristic. Further, we implement the so called “modular discrete Fourier transform” in SageMath [SAGE] for the cyclic group and for the symmetric group $S_{n}$.
The modular discrete Fourier transform for a field $K$ with $\text{char}(K)>0$ uses the isomorphism from the group algebra $K\left[G\right]$ to its block decomposition $\bigoplus_{i}K\left[G\right]e_{i}$, where $1=\sum_{i}e_{i}$ is a decomposition of the identity into primitive orthogonal idempotents. This generalizes the usual Fourier transform which uses Mashke’s theorem to decompose $K\left[G\right]=\bigoplus_{i}\text{End}(V_{i})$ for irreducible representations $V_{i}$ of $G$ so long as $\text{char}(K)\nmid|G|$. Though we do not use them directly, the dimensions of simple modules $D^{\lambda}$ for $p$-regular partitions $\lambda$ are computed for degree $n\leq 6$ in [dim-simple-modules] via the Gram matrix of the bilinear form in the polytabloid basis.
The DFT generalizes readily to the realm of quantum computing. We view a collection of $n$ qubits $c_{0}|0\rangle+c_{1}|1\rangle\in\mathbb{C}^{2}$ which are concatenated in space to correspond to the Hilbert space $\mathcal{H}\cong\left(\mathbb{C}|0\rangle+\mathbb{C}|1\rangle\right)^{\otimes n}\cong\mathbb{C}^{2^{n}}$. Then the Fourier transform is a unitary matrix $U:\mathcal{H}\rightarrow\mathcal{H}$, where we label the eigenstates using binary notation, e.g. for $n=3$, $|0\rangle\otimes|1\rangle\otimes|1\rangle=|011\rangle$ can be written as $|3\rangle$ in the familiar decimal, and we interpret the label modulo $N=2^{n}$. We can then define maps such as $|x\rangle\mapsto|a^{x}(\text{mod}\,N)\rangle$ which is useful in applications like Shor’s algorithm [PS94], where we use the inverse quantum Fourier transform to extract the periodicity of that modular exponentiation viewed as a function $\mathbb{Z}/N\mathbb{Z}\rightarrow\mathbb{Z}/N\mathbb{Z}$. This technique is referred to as “Fourier sampling.”
@@ Acknowledgements
I would like to thank Ezra Brown, Robert Pollack, Sridhar Ramesh, Steve Rosenberg, Travis Scrimshaw, and Jared Weinstein for helpful comments and conversation.
@@ 2. The Discrete Fourier Transform
It is important that the base field, $\mathbb{C}$, is of characteristic zero. This means that we can always invert the order of the cyclic group, $N$, to define the inverse Fourier transform. However, we have thus far suppressed the representation theory and harmonic analysis. Note that $\mathbb{Z}/N\mathbb{Z}$ is an abelian group, so its representations are one dimensional. So we are looking at maps $\chi:\mathbb{Z}/N\mathbb{Z}\rightarrow\mathbb{C}$. This means that $\chi(N)=\chi(1)^{N}=1$, which means $\chi(1)$ is an $N^{th}$ root of unity. Since these are one dimensional, they are irreducible. Thus, there are $N$ representations given by $\chi_{k}(n)=\exp(\frac{-2\pi ikn}{N})$.
@@
Note that $\chi_{k}$ are functions on the group $\mathbb{Z}/N\mathbb{Z}$. Thus, for each group element $n\in\{0,\ldots,N-1\}$ we have a complex number $\chi_{k}(n)\in\mathbb{C}$. Recall that the space of class functions on a group form a vector space. Since the group is abelian, each conjugacy class has one element, so $\chi_{k}$ may be viewed as a vector in $V^{*}\cong\mathbb{C}^{N}$. Separately, we have the group alegbra $V=\mathbb{C}\left[\mathbb{Z}/N\mathbb{Z}\right]\cong\mathbb{C}^{N}$ as a vector space. We can multiply these “vectors” by viewing group elements as a basis and extending the multiplication as we do with polynomials. Label the irreducible representations $V_{k}\cong\mathbb{C}$. Each element $n\in\mathbb{Z}/N\mathbb{Z}$ may be viewed as an endomorphism $n:V\rightarrow V$ by multiplication. That is, if $s=\sum_{j}c_{j}j\in V$ then $n\cdot s=\sum_{k}c_{j}(n\cdot j)$ where the product $n\cdot j=n+j\mod N$ is in $\mathbb{Z}/N\mathbb{Z}$. Therefore we have a linear action of $\mathbb{Z}/N\mathbb{Z}$ on $V$, which is invertible since $-n$ is the multiplicative inverse of $n$. Thus, each $n\in GL(V)$, so this is a representation of the group. $V$ can be decomposed into irreducible representations $V\cong\bigoplus_{k}V_{k}$ and the action of $n$ diagonalized as $n_{k}\in End(V_{k})$. We thus have the decomposition $V=\mathbb{C}\left[\mathbb{Z}/N\mathbb{Z}\right]\cong\bigoplus_{k}\text{End}(V_{k})\cong\bigoplus_{k}\mathbb{C}\chi_{k}=V^{*}$.
@@
So how does the Fourier transform look in this interpretation? Consider the element $x=\left(x_{1},\ldots,x_{N}\right)\in\mathbb{C}^{N}$. We map it to $\hat{x}_{k}=\sum_{n=0}^{N-1}x_{n}\exp\left(\frac{-2\pi ikn}{N}\right)$.
Let us rewrite this formula a bit.
$\hat{x}_{k}=\sum_{n=0}^{N-1}x_{n}\exp\left(\frac{-2\pi k}{N}\right)^{n}=\sum_{n=0}^{N-1}x_{n}\chi_{k}(n)$
@@
We see that our “Fourier coefficients” are just sums of the components of $x$ against the values of the irreducible characters. We obtain $N$ coefficients, and this is exactly the discrete Fourier transform for the group $\mathbb{Z}/N\mathbb{Z}$. This formula has the nice property of being invertible, with inverse given by $x_{n}=\frac{1}{N}\sum_{k=0}^{N-1}\hat{x}_{k}\chi_{k}(-n)$, whose proof uses the orthogonality of characters.
@@
This is a bit of a special case, in which the number of representations equals the order of the group, i.e. every representation is degree 1. This only occurs for abelian groups in algebraically closed fields of characteristic 0. For nonabelian groups, we need to use matrices.
@@ 3. Fourier Transform for Non-Abelian Groups
Let $G$ be a finite group and let $\widehat{G}$ denote the set of irreducible representations of $G$, also a finite set. Over a field of characteristic zero, we have Maschke’s theorem:
@@
This theorem tells us that as a $K\left[G\right]$-module we may decompose $K\left[G\right]$ into a direct sum of simple $K\left[G\right]$-modules. That is, $K\left[G\right]=\bigoplus_{i}M_{i}$. It turns out these simple $M_{i}$ are just the endomorphism algebras of irreducible representations, i.e. $K\left[G\right]=\bigoplus_{i}End(V_{i})$. Thus, an element of the group algebra may be viewed as a “block diagonal” matrix acting on itself, with each block corresponding to an irreducible representation. This is a special case of Wedderburn’s theorem.
@@
where $f(a)\in K$ and each $\rho(a)$ is a $d_{\rho}\times d_{\rho}$ matrix. This formula is invertible yielding
$f(a)=\frac{1}{|G|}\sum_{g\in\hat{G}}d_{\rho}\text{Tr}(\rho(a^{-1})\hat{f}(\rho)).$
It is clear we require $\text{char}(K)\nmid|G|$, both in order to comply with the hypothesis of Mashke’s Theorem, and to avoid division by zero. We need to know the dimension numbers of irreducible representations, $d_{\rho}$. We are dealing with matrices, so we are taking the matrix trace now in the sum.
@@ 4. The Symmetric Group over $\mathbb{C}$
For the symmetric group over $\mathbb{C}$, Maschke’s theorem applies. The irreducible representations are labeled by Young tableaux $\lambda$ corresponding to partitions of $m$. A tabloid is an equivalence class of Young tableaux where two labelings are equivalent if one is obtained from the other by permuting the entries of each row. For each Young tableau $T$ of shape $\lambda$ let $\{T\}$ be the corresponding tabloid. For each $\lambda$, we obtain a simple modules $V^{\lambda}$ with basis given by
@@
where $Q_{T}$ is the subgroup of permutations, preserving (as sets) all columns of $T$ and $\epsilon(\sigma)$ is the sign of the permutation $\sigma$. The Specht module of the partition $\lambda$ is the module generated by the elements $E_{T}$ as $T$ runs through all tableaux of shape $\lambda$. The dimension of the Specht module $V^{\lambda}$ is the number of standard Young tableaux of shape $\lambda$. It is given by the hook length formula. Over fields of characteristic 0 the Specht modules are irreducible and form a complete set of irreducible representations.
We now have everything we need to compute the non-abelian Fourier transform, which is implemented in [DFT] for the symmetric group (though it generalizes to any finite group given a labeling for the representations).
@@
There is a well-known isomorphism of categories between $K\left[G\right]$-modules are representations of the group $G$ over the field $K$. The “regular representation” where $k\left[G\right]$ acts on itself by multiplication decomposes into a sum of irreducible representations. When the characteristic of the field $p$ does not divide the order of the group $|G|$ we can use Maschke’s theorem as stated above. The key element of the proof is the construction of an “average” or invariant inner product, $\langle x,y\rangle_{G}=\frac{1}{|G|}\sum_{g\in G}\langle gx,gy\rangle$, where juxtaposition denotes the group action in a representation. Note that typically we require $K$ be algebraically closed, since otherwise a more complicated statement holds: $K\left[G\right]$ is a product of matrix algebras over division rings over $K$. In the case of finite fields, any division ring over $F_{p}$ must be a field by Wedderburn’s little theorem, and the only finite fields of characteristic $p$ are the extensions $F_{q}$ where $q=p^{\nu}$.
When $p\mid|G|$, the situation is more complicated and is the subject of modular representation theory. In our case, we are primarily concerned with cyclic groups and the symmetric group $S_{m}$. In general, we no longer have Maschke’s theorem giving us a decomposition of the group algebra into simple modules. Instead, we have a decomposition into “blocks”,
@@
where $\sum_{i}e_{i}=1$ is a decomposition of the identity into central primitive orthogonal idempotents. In certain cases these idempotents are readily computable.
@@ 6. The Discrete Fourier Transform over a Ring
We implement this procedure in SageMath code. [SAGE] [DFT]
@@
When $R=\mathbb{F}_{p}$, the multiplicative group $\mathbb{F}_{p}^{\times}\cong\mathbb{Z}/(N-1)\mathbb{Z}$ is cyclic. By Fermat’s little theorem, $\alpha^{p-1}=1$. Thus, we require $N|(p-1)$. This implies that $p>N+1$, which means that the characteristic we choose must be bigger than the size of the list we are transforming. This is quite a constraint!
@@ 7. Cyclic Groups over $\mathbb{F}_{p}$
The question of how to classify all finite dimensional representations of $\mathbb{Z}/N\mathbb{Z}$ over an arbitrary field $F$ can be studied using the structure theorem for finitely-generated modules over a principal ideal domain, in this case $F\left[x\right]$. The structure theorem asserts that any finitely-generated module is uniquely a finite direct sum of modules of the form $F\left[x\right]/p(x)^{r}$ where $p\in F\left[x\right]$ is irreducible and $r$ is a non-negative integer.
If $T$ is an operator acting on $F^{k}$ for some $n$, then $F^{k}$ becomes a finitely-generated module over $F\left[x\right]$ with $x$ acting by $T$. This gives a representation of the cyclic group $\mathbb{Z}/N\mathbb{Z}$ if and only if $T^{N}=1$, in which case the summands $F\left[x\right]/q(x)^{r}$ in the decomposition of $F^{k}$ must have the property that $q(x)^{r}|(x^{N}-1)$.